OpenStudy (anonymous):

how can I get the Value of an exponent ? Examples 3^x= 2187*81

ganeshie8 (ganeshie8):

see if 2187 can be written as powers of 3

OpenStudy (akashdeepdeb):

3^x = 2187*81 Now break up the RHS in the form of exponents with base 3 3^x = 243*9*81 3^x = 3^5 * 3^2 * 3^4 \[3^x = 3^{5+2+4}\] [Basic laws of exponents] So \[3^x = 3^{11}\] Now if the bases are same ie. 3 then even their powers have to be the same. Understood? :)

OpenStudy (akashdeepdeb):

@EinsteinMorse ?

OpenStudy (anonymous):

Yeah thanks . Can you give some sort of examples?

OpenStudy (anonymous):

You need two identities to solve this: \[x ^{n} = n*\log(x)\] and \[a*b = \log(a) + \log(b)\] Take into consideration the info AkashdeepDeb provided, and it becomes straight forward algebra. (working on detailed steps...)